NERSCPowering Scientific Discovery Since 1974

January 3

Attending:  Eric Iwona Mike Larry JeffP

Cluster Utilization:  Filled to capacity for the past three weeks now.  Down to 1700 cores now because a few cores were taken for other purposes.

Recent Outages/Incidents:  Things were quiet over the break.  Backups have been finishing which is indicative of too file system performance.  An additional metadata was added so the home and common no longer share.  Efforts are underway to get users doing data transfers to use pdsfdtn1 and pdsfdtn2.

Upcoming Downtimes:  Nothing planned.

Procurements:  Nothing underway but will do one before April.

SL4 Retirement:  Still tentatively scheduled for 3/1/12.

Project Accounts:  Nothing to report.

Grid Issues:  Alice problems with Alien due to security blocking connections have been resolved.

Topics from Users:

Mike:  Still planning to work on ATLAS storage reconfiguration - need to decide when to do it.

Iwona:  SL6 testing is underway.